OpenStudy (anonymous):

If f(x) = 3x - 1 and g(x) = x2 - 2x - 2, what is f(g(x))

OpenStudy (anonymous):

Fog(x)=3(x^2-2x-2)-1 =3x^2-6x-6-1 =3x^2-6x-7

OpenStudy (experimentx):

\( f(g(x)) = f(x^2 - 2x - 2) = 3(x^2 - 2x - 2) - 1 \) <--- simplify it.

OpenStudy (anonymous):

put g(x)in place of x in f(x)